Related Rates: Problem Solving


Direction: Solve each word problem involving a related rates. Make a sketch of your work

if necessary, and then apply differentiation.


An automatic travelling at a rate of 30 feet rate second is approaching an

intersection. When the automobile is 120 feet from the intersection, a truck

traveling at a rate of 40 feet per second crosses the intersection. The automobile

and the truck going, separating 2 seconds after the truck leaves the intersection?


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/03/2022 11 : 00 PM


1
Expert's answer
2022-05-05T14:27:39-0400

truck goes for 2 seconds 40x2=80 feet

Automobile goes for 2 seconds 30x2=60 feets.

120-60=60 feets automobile before intersection.

So for 2 seconds distance between truck and automobile is "d=\sqrt{60^2+80^2}=\sqrt{10000}=100 feets"